1-the rth term in a sequence is 2+3r. Find the first three terms of the sequence.

Answer by mathsnight(4)

Given,
the rth term, ar= 2+3r, thus, the first 3 terms are: a1 = 2+3(1)=5,
a2 = 2+3(2)=8,
a3 = 2+3(3)=11.. Thus the first 3 terms are 5,8 and 11. This is A.P with d=3.